Analyze the annual reports of Flour Mills of Fiji (2020), RB Patel Ltd (2020) and Paradise Beverages Ltd (2019) from South Pacific Stock Exchange (SPSE) and answer the following questions. 

 

1.     Calculate the share value using Dividend Discount Model, assuming dividend will grow 3% for Flour Mills of Fiji, 6% for RB Patel Ltd and 9% for Paradise Beverages constantly for the companies on their respective base year. 


1
Expert's answer
2021-04-20T17:24:24-0400

Let the discount rate be 10, the dividends are 5,6,7, respectively, then we will find the price

"P=\frac{D}{r-g}"


"P1=\frac{5}{0.1--0.03}=71,43"

"P2=\frac{6}{0.1--0.06}=150"

"P1=\frac{7}{0.1--0.09}=700"